### v2

@@ Appendix F Confounding function
F.1 Variance of confounding function
Suppose we now want to compute the variance of the confounding function evaluated at $\mathbf{x}_{*}$. This is given by $\eta(\mathbf{x}_{*})=f^{\text{e}}(\mathbf{x}_{*})-f^{o}(\mathbf{x}_{*})$, or in matrix notation $\eta(\mathbf{x}_{*})=\begin{pmatrix}1\\ -1\end{pmatrix}^{T}\mathbf{f}(\mathbf{x}_{*})$. By standard properties of the multivariate Gaussian distribution, it follows that
@@ F.1 Variance of confounding function
$V_{\eta}(\mathbf{x}_{*})$ $=\begin{pmatrix}1\\ -1\end{pmatrix}^{T}K_{*}\left(\mathbf{x}_{*},\mathbf{x}_{*}\right)\begin{pmatrix}1\\ -1\end{pmatrix}$
@@
Appendix G Multivariate simulation studies
G.1 Multivariate Data Generation Process
@@
Simulation 1: Noninear Potential Outcomes, linear confounding, linear CATE
@@
Simulation 2: Noninear Potential Outcomes, nonlinear confounding, nonlinear CATE
@@ G.2 Multivariate Simulations: Results
In the second simulation setting, where both the CATE and the confounding exhibit more complex, non-linear structures, the performance of models relying on correct specification deteriorates. In contrast, Causal-ICM maintains robust performance, being outperformed only by the Integrative HTE. Importantly, this superior performance of the Integrative HTE is a consequence of having a correctly specified outcome model; in realistic scenarios where model specification is uncertain or misspecified, we expect Causal-ICM to offer more reliable and consistent results.
Appendix H Extrapolation with Causal-ICM
@@
Appendix I Sensitivity to the parameter $\rho$, the sample size of the observational study, the kernel choice and the degree of overlap
@@
Sample size Mean RMSE (SD)
$n^{(e)}=100,\;n^{(0)}=100$ 0.879 (0.484)
$n^{(e)}=100,\;n^{(0)}=500$ 0.797 (0.524)
$n^{(e)}=100,\;n^{(0)}=1000$ 0.731 (0.448)
$n^{(e)}=100,\;n^{(0)}=2000$ 0.793 (0.481)
Table 4: Mean RMSE for different sample sizes, Univariate simulation 2; Results obtained over 50 different simulated datasets.
@@
Table 5: Mean RMSE for different degrees of overlap between the RCT and observational study for the second simualation setting of the main text (Univariate simulation 2); The degree of overlap is controlled through the coefficients of the model for study participation; Results obtained over 100 different simulated datasets.
The results above show that
  1. 1.
    $\rho$ indeed plays an important role in controlling the influence of the unobserved confounding
  2. 2.
    Causal-ICM retains strong performance for different kernels, or even as the sample size of the (confounded) observational data grows
  3. 3.
    The Causal-ICM performs well even with low overlap between the two studies, although optimal results are obtained with full overlap.
Appendix J Priors for latent processes, extentions to non-Gaussian likelihoods, and computational complexity
In the current implementation of Causal-ICM, both $u_{1}(x)$ and $u_{2}(x)$ share the same kernel hyperparameters (lengthscale, variance). However, extensions to the model do exist that allow $u_{1}(x)$ and $u_{2}(x)$ to originate from distinct Gaussian Processes, such as in the Linear Model of Coregionalization (LCM) Alvarez et al. (2012). Exploring these extensions represents a promising direction for future research.
With respect to extensions to non-Gaussian likelihoods, our theoretical results currently rely on the conjugacy of the Gaussian likelihood, which greatly simplifies the proofs. Nevertheless, we anticipate that similar results can be extended to other likelihoods that have closed-form posterior variances, such as the Student-T process. We expect these extensions to be non-trivial, however, and so leave these for future work.
For likelihoods requiring approximation methods, such as the Bernoulli likelihood, the analysis becomes significantly more intricate compared to conjugate cases. Nevertheless, we are actively exploring this direction as part of our efforts to extend causal-ICM to accommodate binary outcomes. A promising avenue involves employing a sigmoid link function to transform Gaussian Process outputs into the [0,1] interval, enabling compatibility with binary targets. While we expect comparable results in this setting, posterior approximations will be necessary. This extension represents a significant focus of our ongoing research and will play a central role in our future developments.
Last, GPs are known for their computational complexity. In our examples, we utilize the full GP as the datasets are not too large. However, their cubic computational scaling with the number of data points is indeed a well-recognized limitation when datasets are larger. To mitigate this issue, we suggest using sparse GPs as an approximation method, where we reduce computational cost by relying on a subset of the data (i.e. inducing points) to approximate the full GP model.

### v3

@@ Appendix G Confounding function
G.1 Variance of confounding function
Suppose we now want to compute the variance of the confounding function evaluated at $\mathbf{x}_{*}$. This is given by $\eta(\mathbf{x}_{*})=f^{\text{e}}(\mathbf{x}_{*})-f^{o}(\mathbf{x}_{*})$, or in matrix notation $\eta(\mathbf{x}_{*})=\begin{pmatrix}\phantom{-}1\\ -1\end{pmatrix}^{T}\mathbf{f}(\mathbf{x}_{*})$. By standard properties of the multivariate Gaussian distribution, it follows that
@@ G.1 Variance of confounding function
$V_{\eta}(\mathbf{x}_{*})$ $=\begin{pmatrix}\phantom{-}1\\ -1\end{pmatrix}^{T}K_{*}\left(\mathbf{x}_{*},\mathbf{x}_{*}\right)\begin{pmatrix}\phantom{-}1\\ -1\end{pmatrix}$
@@
Appendix H Multivariate simulation studies
@@
Simulation 1: Nonlinear Potential Outcomes, linear confounding, linear CATE
@@
Simulation 2: Nonlinear Potential Outcomes, nonlinear confounding, nonlinear CATE
@@
In the second simulation setting, where both the CATE and the confounding exhibit more complex, non-linear structures, the performance of models relying on correct specification deteriorates. In contrast, Causal-ICM maintains robust performance, being outperformed only by the Integrative HTE. Importantly, this superior performance of the Integrative HTE is a consequence of having a correctly specified outcome model; In the multivariate simulation settings, parametric competitors can outperform Causal-ICM when their outcome and confounding models are correctly specified. Table 2 shows, however, that under misspecification Causal-ICM becomes more robust and substantially outperforms Integrative HTE in the second multivariate simulation setting.
Table 2: RMSE in the second multivariate simulation setting under model misspecification.
Model Mean RMSE (SD)
GP (exp) 3.422 (2.186)
GP (obs) 3.580 (0.796)
Experimental Grounding 3.582 (1.713)
Integrative HTE 21.883 (3.330)
Causal-ICM 2.586 (1.513)
Appendix I Extrapolation with Causal-ICM
@@
Appendix J Sensitivity to $\rho$, Observational Sample Size, Kernel Choice, and Degree of Overlap; and Performance Within and Outside the RCT Support
@@
Table 5: Performance (RMSE) under increasing observational sample size in simulation setting 2. Results are based on 50 simulated datasets.
Model Observational sample size Mean RMSE (SD)
Causal-ICM (ours) 200 1.078 (0.810)
Integrative HTE 200 1.080 (0.441)
Power Likelihood 200 1.023 (0.559)
Experimental Grounding 200 1.691 (0.597)
GP Observational 200 1.512 (0.226)
Causal-ICM (ours) 500 1.071 (0.652)
Integrative HTE 500 1.406 (0.461)
Power Likelihood 500 1.231 (0.667)
Experimental Grounding 500 1.646 (0.547)
GP Observational 500 1.475 (0.146)
Causal-ICM (ours) 1000 0.929 (0.695)
Integrative HTE 1000 1.546 (0.303)
Power Likelihood 1000 0.975 (0.506)
Experimental Grounding 1000 1.652 (0.452)
GP Observational 1000 1.414 (0.128)
Causal-ICM (ours) 2000 0.934 (0.674)
Integrative HTE 2000 1.576 (0.318)
Power Likelihood 2000 1.161 (0.606)
Experimental Grounding 2000 1.660 (0.443)
GP Observational 2000 1.377 (0.062)
@@
Table 6: Mean RMSE for different degrees of overlap between the RCT and observational study for the second simulation setting of the main text (Univariate simulation 2); The degree of overlap is controlled through the coefficients of the model for study participation; Results obtained over 100 different simulated datasets.
The results above highlight the importance of $\rho$ in controlling the influence of unobserved confounding: as $\rho$ increases from 0 to 0.8, the RMSE decreases steadily, reflecting a stronger borrowing of information from the observational study, before rising again at $\rho=1.0$ where the confounding is no longer accounted for. Across different kernel choices, Causal-ICM retains strong and consistent performance, with the RBF kernel achieving the lowest mean RMSE, suggesting some robustness to this modelling choice, though the differences are moderate.
Table 5 reports results under increasing observational sample size while keeping the experimental sample size fixed. Causal-ICM remains competitive as the observational sample grows substantially larger than the RCT, and in fact improves with larger observational samples, in contrast to Integrative HTE whose performance degrades noticeably. The performance of the experimental grounding method of Kallus et al. (2018), the power likelihood of Lin et al. (2025), and the T-learner trained on observational data does not seem to deteriorate with the increasing observational sample size, but it remains inferior to our method. Finally, Table 6 shows that Causal-ICM performs well even under low overlap between the two studies, with optimal results obtained under full overlap, as expected. Taken together, these results suggest that Causal-ICM is robust across a range of practical conditions that may arise when combining experimental and observational data.
To complement the aggregate RMSE results in the main text, Table 7 reports mean squared error separately inside and outside the support of the RCT for simulation setting 2.
Table 7: Mean in-sample and out-of-sample MSE in simulation setting 2. In-sample refers to the support of the randomized trial; out-of-sample refers to regions outside the trial support.
Method Mean in-sample MSE (SD) Mean out-of-sample MSE (SD)
GP exp 15.302 (15.595) 9.645 (13.259)
GP obs 343.387 (436.318) 310.823 (412.965)
Kallus GP 3.048 (3.141) 43.461 (42.069)
Integrative HTE 0.307 (0.114) 3.153 (0.919)
Causal-ICM 0.252 (0.099) 1.071 (0.722)
Causal-ICM achieves the lowest in-sample MSE among all methods, and most strikingly, retains strong performance outside the RCT support, where competing methods deteriorate considerably. In particular, Kallus GP degrades sharply out-of-sample despite performing reasonably well within the trial support, suggesting it does not generalise well beyond the experimental region. Causal-ICM, by contrast, benefits from the broader covariate coverage of the observational study to extrapolate more reliably, underscoring one of the key practical advantages of integrating the two data sources.
Appendix K Runtime results
Training multi-task Gaussian processes scales cubically in the total sample size, although the intrinsic coregionalization structure can often be exploited to improve efficiency in practice. For larger datasets, sparse GP approximations with inducing points provide a natural route to scalability.
Table 8: Runtime performance in seconds for simulation setting 2, based on 100 datasets.
Method Mean (s) SD (s) Median (s) Min (s) Max (s)
GP exp 0.244369 0.114056 0.236458 0.089087 0.489595
GP obs 2.232287 0.639261 2.077287 1.417782 6.207515
Experimental Grounding GP 2.394518 0.951331 2.181754 1.566869 10.213725
Causal-ICM 4.559152 0.782487 4.513439 3.310538 8.185129
Integrative HTE 6.006100 0.416126 5.964500 5.222000 7.494000